# OccamLLM: Fast and Exact Language Model Arithmetic in a Single Step

### v1

@@ B.1 Alternative Architectures
Regarding transformers and RNNs, we believe that OccamNet possesses a key advantage of being interpretable; by looking at the weights, it is possible for a human to see what the architecture is doing. We beleive that this interpretability will also make it easier for a decoder to initialize properly, as opposed to an RNN or transformer where the relation between the weights and the probability distribution is substantially more complex. This leads us to our key point. Transformers and RNNs are effective for modeling complex multimodal distributions, but for this problem, we want to select a single function for each token, so the extra expressivity of these models is unneeded and likely detrimental to performance. We beleive that OccamNet, being a much simpler architecture, enables better parameter efficiency and performance.
B.2 Alternative Losses
@@
We considered two loss functions which are natural when optimizing a probability distribution: 1) a cross entropy loss, and 2) a reinforcement learning loss.
Cross entropy losses are effective, but they require access to the ground-truth distribution. As such, this approach may not scale well for OccamNet.
Reinforcement learning losses, on the other hand are effective for exploring large search spaces. We find that using REINFORCE with a delta function for our reward is most effective because it mimics a cross entropy loss. We rescale the loss by the total reward for all the samples to ensure that correct functions sampled only a few times still receive a large training step update.
Appendix C Background on OccamNet
Much of this section is modified from [11].
@@ Appendix C Background on OccamNet
    In Section C.1, we describe OccamNet’s architecture in more detail.
@@
    In Section C.2, we describe OccamNet’s sampling process.
@@
    In Section C.3, we describe OccamNet’s probability distribution.
@@
    In Section C.4, we describe OccamNet’s initialization process.
C.1 OccamNet Architecture
As described in the main text, we start from a predefined collection of $N$ primitive functions $\mathcal{P}$. OccamNet represents a distribution over compositions of functions in $\mathcal{P}.$ From now on, we denote the $i$th primitive function in the $l$th layer as $\phi_{i}^{(l)}$. We 0 index the primitives but 1 index the layers, because we treat the inputs as the 0th layer. So, for example, in Figure 2a, $\phi_{2}^{(1)}=\phi_{2}^{(2)}=\phi_{2}^{(3)}=\mathsf{sin}$.
Figure 5: A demonstration of the dropped connections from sampled paths in OccamNet. All light grey paths are dropped from the final symbolic form of the sampled function because they are not directly connected to the outputs. These paths are unnecessarily computed during OccamNet’s training process, leading to potential slowdowns in training.
Each OccamNet layer consists of two sublayers, which we denote the arguments and image sublayers, shown in Figure 5a. For an $L$-layer OccamNet, each of these sublayers is reproduced $L$ times. The $l$th softmax layer connects the $(l-1)$th image layer with the $l$th arguments layer. For $1\leq l\leq L$, we denote the $l$th arguments sublayer hidden state as $\widetilde{\mathbf{h}}^{(l)}$ and the $l$th image sublayer hidden state as $\mathbf{h}^{(l)}$. So, $\widetilde{\mathbf{h}}^{(2)}$ would represent the We further write
$\mathbf{\widetilde{h}}^{(l)}=\left[\widetilde{h}^{(l)}_{1},\dots,\widetilde{h}^{(l)}_{M^{(l)}}\right],\;\;\mathbf{h}^{(l)}=\left[h^{(l)}_{1},\dots,h^{(l)}_{N^{(l)}}\right],$ (4)
@@ C.1 OccamNet Architecture
$M^{(l)}=\sum_{0\leq k\leq N^{(l)}}\alpha\left[\phi_{k}^{(l)}\right]$
and $\alpha[\phi]$ is the arity of function $\phi.$ We also define $\mathbf{h}^{(0)}$ to be the input layer (an image sublayer) and $\mathbf{\widetilde{h}}^{(L+1)}$ to be the output layer (an arguments sublayer).
In a standard OccamNet layer, each primitive is repeated exactly once in each layer. However, in Complete OccamNet, each primitive in the $l$th layer is repeated $A^{L-l}$ times, where $A$ is the maximum arity of the primitives. This is shown in Figure 6 in the transition from 6a) to 6b). Complete OccamNet also concatenates each image layer to the next image layer, as shown in Figure 6c.
C.2 Sampling from OccamNet
In this section, we more carefully describe OccamNet’s sampling process. For each softmax layer, we sample a connection to each arguments layer node from the softmax of the weights. In particular, if $\mathbf{w}_{i}^{(l)}$ are the weights of the $l$th softmax layer leading to the $i$th node of the $l$th argument’s layer, when we sample we produce a sparse matrix
@@ C.2 Sampling from OccamNet
where the $\mathsf{SAMPLE}$ function samples a one-hot row vector for each row based on the categorical probability distribution defined by $\mathsf{softmax}(\mathbf{w})$. To evaluate a forward pass through this sample, we simply evaluate a forward pass through the network, treating the sampled sparse matrices from the softmax layers as the weights of linear layers:
@@
$h^{(l)}_{i}=\phi_{i}^{(l)}\left(\widetilde{h}_{j}^{(l)},\dots,\widetilde{h}_{j+\alpha(\phi_{i})-1}^{(l)}\right),\;\;j=\sum_{0\leq k<i}\alpha(\phi_{k}^{(l)}).$ (7)
Figure 6: Skip connections. Nodes are color coded with lines indicating the origin of the reused neurons.
C.3 OccamNet’s Probability Distribution
OccamNet parametrizes a probability distribution over all functions which it can sample. In particular, when OccamNet samples a function, it is really sampling a directed acyclic graph (DAG) which defines a computational path to compute a function. The probability of sampling a function is equal to the product of the probabilities of the connections in the DAG which are connected to the output node.
@@ C.3 OccamNet’s Probability Distribution
C.4 Initialization
This section describes how we calculate $\mathbf{W}_{i}^{(0)}$ from the main text.
The algorithm to initialize all functions with equal probability establishes as an invariant that, after assigning the weights up to the $l$th layer, all paths leading to a given node in the $l$th argument layer have equal probabilities. Then, each argument layer node has a unique corresponding probability, the probability of all paths up to that node. We denote the probability of the $i$th node in the $l$th argument sublayer as $\widetilde{p}_{i}^{(l)},$ because it is the probability of any path leading to the $i$th node in the $l$th argument sublayer. Because each argument layer node has a corresponding probability, each image layer node must also have a unique corresponding probability, which, for the $i$th node in the $l$th image sublayer, we denote as $p_{i}^{(l)}.$ Again, we use the notation $p_{i}^{(l)}$ because this is the probability of any path leading to the $i$th node in the $l$th image sublayer. These image layer probabilities are given by
[…]
Our algorithm starts with input layer, or the 0th image layer. Paths to any node in the input layer have no edges so they all have probability 1. Thus, we initialize $p_{i}^{(0)}=1$ for all $i$. As the algorithm iterates through all subsequent $T$-Softmax layers, the invariant established above provides a system of linear equations involving the desired connection probabilities, which the algorithm solves. The algorithm groups the previous image layer according to the node probabilities, obtaining a set of ordered pairs $\{({p^{\prime}}_{a}^{(l)},n_{a}^{(l)})\}_{i=a}^{k}$ representing $n_{a}^{(l)}$ nodes with probability ${p^{\prime}}_{a}^{(l)}$ in the $l$th layer. Note that if two image nodes have the same probability $p_{i}^{(l)}=p_{j}^{(l)}$, then the edges between any argument node in the next layer and the two image nodes must have the same probability in order to satisfy the algorithm’s invariant: $p_{i}^{(l,k)}=p_{j}^{(l,k)}$. Then, we define ${p^{\prime}}_{a}^{(l,i)}$ as the probability of the edges between the image nodes with probability ${p^{\prime}}_{a}^{(l)}$ and the $i$th argument $P$-node of the $l$th layer. The probabilities of the edges to a given $P$-node sum to 1, so for each $j$, we must have $\sum_{a}n_{a}^{(l)}{p^{\prime}}_{a}^{(l,i)}=1.$ Further, the algorithm requires that the probability of a path to a $P$-node through a given connection is the same as the probability of a path to that $P$-node through any other connection. The probability of a path to the $i$th $P$-node through a connection with probability ${p^{\prime}}_{a}^{(l,i)}$ is ${p^{\prime}}_{a}^{(l)}{p^{\prime}}_{a}^{(l,i)},$ so we obtain the equations ${p^{\prime}}_{0}^{(l)}{p^{\prime}}_{0}^{(l,i)}={p^{\prime}}_{a}^{(l)}{p^{\prime}}_{a}^{(l,i)},$ for all $a$ and $i$. These two constraints give the vector equation
$\begin{bmatrix}n_{0}^{(l)}&n_{1}^{(l)}&n_{2}^{(l)}&\cdots&n_{k}^{(l)}\\ {p^{\prime}}_{0}^{(l)}&-{p^{\prime}}_{1}^{(l)}&0&\cdots&0\\ {p^{\prime}}_{0}^{(l)}&0&-{p^{\prime}}_{2}^{(l)}&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ {p^{\prime}}_{0}^{(l)}&0&0&\cdots&-{p^{\prime}}_{k}^{(l)}\end{bmatrix}\begin{bmatrix}{p^{\prime}}_{0}^{(l,j)}\\ {p^{\prime}}_{1}^{(l,j)}\\ {p^{\prime}}_{2}^{(l,j)}\\ \vdots\\ {p^{\prime}}_{k}^{(l,j)}\end{bmatrix}=\begin{bmatrix}1\\ 0\\ 0\\ \vdots\\ 0\end{bmatrix},$ (9)
for all $1\leq j\leq M$. The algorithm then solves for each ${p^{\prime}}_{a}^{(l,j)}.$
After determining the desired probability of each connection of the $l$th layer, the algorithm computes the SPL weights ${\mathbf{w}^{\prime}}^{(l,j)}$ that produce the probabilities ${p^{\prime}}_{a}^{(l,j)}$. Since there are infinitely many possible weights that produce the correct probabilities, the algorithm sets ${w^{\prime}}_{0}^{(l,j)}=0.$ Then, the algorithm uses the softmax definition of the edge probabilities to determine the required value of $\sum_{n=1}^{k}\exp({w^{\prime}}_{n}^{(l,j)}/T^{(l)})$:
${p^{\prime}}_{0}^{(l,j)}$ $=\frac{\exp({w^{\prime}}_{0}^{(l,j)}/T^{(l)})}{\sum_{n=1}^{k}\exp({w^{\prime}}_{n}^{(l,j)}/T^{(l)})}$
$=\frac{1}{\sum_{n=1}^{k}\exp({w^{\prime}}_{n}^{(l,j)}/T^{(l)})}$
so
$\sum_{a=1}^{k}\exp({w^{\prime}}_{a}^{(l,j)}/T)=1/{p^{\prime}}_{0}^{(l,j)}.$
Substituting this equation into the expression for the other probabilities gives
${p^{\prime}}_{a}^{(l,j)}$ $=\exp({w^{\prime}}_{a}^{(l,j)}/T^{(l)})/\left(\sum_{n=1}^{k}\exp({w^{\prime}}_{n}^{(l,j)}/T^{(l)})\right)$
$={p^{\prime}}_{0}^{(l,j)}\exp({w^{\prime}}_{a}^{(l,j)}/T^{(l)}).$
Solving for ${w^{\prime}}_{i}^{(l,j)}$ gives
${w^{\prime}}_{a}^{(l,j)}=T^{(l)}\log\left({p^{\prime}}_{a}^{(l,j)}/{p^{\prime}}_{0}^{(l,j)}\right),$ (10)
which the algorithm uses to compute ${w^{\prime}}_{i}^{(l,j)}$.
After determining the weights ${w^{\prime}}_{a}^{(l,j)}$ the algorithm assigns them to the corresponding $w_{a}^{(l,j)}.$ In particular, if the $i$th image node has probability ${p^{\prime}}_{a}^{(l)},$ the weights of edges to the $i$th node are given by $w_{i}^{(l,j)}={w^{\prime}}_{k}^{(l,j)},$ for all $j.$ The algorithm then determines the values of $\widetilde{p}_{j}^{(l+1)},$ given by $\widetilde{p}_{j}^{(l+1)}=p_{i}^{(l)}p_{i}^{(l,j)}.$ Finally, the algorithm determines $p_{i}^{(l+1)}$ using Equation 8 and repeats the above process for subsequent layers until it reaches the end of the network.

### v2

@@ C.1 Alternative Architectures
Regarding transformers and RNNs, we believe that OccamNet possesses a key advantage of being interpretable; simply by looking at the weights, it is possible for a human to determine which functions OccamNet assigns a high probability. We believe that this interpretability will make OccamNet easy for a decoder to initialize with the desired distribution. On the other hand, an RNN or transformer have substantially more complex relations between the weights and corresponding probability distribution, which we hypothesize would make learning a decoder for such models difficult.
This leads us to a key point: transformers and RNNs are effective for modeling complex multimodal distributions, but for this problem, we want to select a single function for each token, so the extra expressivity of these models is unneeded and likely detrimental to performance. We believe that OccamNet, a much simpler architecture, enables better parameter efficiency and performance.
C.2 Alternative Losses
@@
We considered two loss functions which are natural when optimizing a probability distribution: 1) a cross-entropy loss, and 2) a REINFORCE [34] loss. Each of these requires only a slight modification to reach Equation 3. This discussion thus illustrates how our loss combines benefits from both the cross-entropy and the reinforcement-learning losses.
Cross-Entropy Loss
The cross-entropy loss is effective at modeling probability distributions. Given a ground truth distribution $q_{x}[f]$ conditioned on the input text $x$, the cross-entropy loss is given by
$\mathcal{L}(x,y;W)=-\sum_{f}q_{x}[f]\log p_{W}[f].$ (4)
Unfortunately, for OccamLLM, the ground-truth distribution $q_{x}[f]$ is not uniquely specified. In particular the only constraints on $q_{x}[f]$ are that it is is normalized and satisfies $q_{x}[f]=0$ if $f$ is not the desired function (i.e., $f(x)\neq y$). Since the same function can be represented in many ways in the OccamNet network (a property true of many function representations), multiple $f$ may satisfy $f(x)=y$, so $q_{x}$ is underparametrized.
The most natural choice for $q_{x}$ is to weight each valid function equally:
$q_{x}[f]=\begin{cases}c_{x}&\text{if }f(x)=y\\ 0&\text{otherwise}\end{cases}$ (5)
where $c_{x}$ is a constant chosen such that $q_{x}$ is normalized, given by the inverse of the number of functions $f$ satisfying $f(x)=y$. However, determining $c_{x}$ requires testing every possible function $f$, which may be infeasible for large OccamNet networks. Further, this $q_{x}$ requires OccamNet to learn a superposition of functions, which may be challenging given its relatively low parameter count.
Another option is to choose a canonical form $f^{*}$ for each function and to set $q_{x}$ to be a 1-hot distribution that is nonzero only at $f^{*}.$ Although this removes the challenge of learning a superposition, it still requires sampling nearly all functions in OccamNet due to the sparsity of $q_{x}.$
Ideally, we would like to find a $q_{x}$ with the following conditions:
  • It enables the cross-entropy loss to be calculated by sampling from OccamNet. This allows us to avoid needing to iterate through and evaluate every $f(x)$ each time we compute the loss, since we can instead obtain a Monte-Carlo estimate.
  • It is minimized when $p_{W}$ is a 1-hot probability distribution. This ensures that OccamNet can represent the optimal distribution.
  • It has $q_{x}[f]\neq 0$ for all $f$ satisfying $f(x)=y$. This improves sample-efficiency by increasing the probability of sampling an $f$ with $q_{x}[f]>0$.
A solution is to set
$q_{x}[f]=\begin{cases}c_{x}p_{W}[f]&\text{if }f(x)=y\\ 0&\text{otherwise}\end{cases}$ (6)
where $c_{x}$ is chosen such that $q_{x}$ is normalized. This gives a loss
$\mathcal{L}(x,y;W)$ $=-\sum_{f}q_{x}[f]\cdot\log p_{W}[f]$
$=-\sum_{f}c_{x}p_{W}[f]\cdot\delta(f(x)-y)\cdot\log p_{W}[f]$
$\approx-\frac{c_{x}}{N}\sum_{f\sim p_{W}}\delta(f(x)-y)\cdot\log p_{W}[f]$
$\approx-\frac{\sum_{f\sim p_{W}}\delta(f(x)-y)\cdot\log p_{W}[f]}{\sum_{f\sim p_{W}}\delta(f(x)-y)},$
where
$\delta(f(x)-y)=\begin{cases}1&\text{if }f(x)=y\\ 0&\text{otherwise}\end{cases}$ (7)
and in the last step we used the fact that $c_{x}$ can be approximated as
$c_{x}=\frac{1}{\sum_{f}p_{W}[f]\delta(f(x)-y)}\approx\frac{N}{\sum_{f\sim p_{W}}\delta(f(x)-y)}.$
This loss is easily computed by sampling from $p_{W}$, it satisfies $q_{W}>0$ for all $f$ satisfying $f(x)=y,$ and it is minimized when $p_{W}$ is a delta function centered at any $f$ satisfying $f(x)=y,$ as desired.
Note that
$\mathcal{L}(x,y;W)=-\frac{\sum_{f\sim p_{W}}\delta(f(x)-y)\cdot\log p_{W}[f]}{\sum_{f\sim p_{W}}\delta(f(x)-y)}$ (8)
is exactly the loss given in Equation 3 with $R(f(x),y)=\delta(f(x)-y)$. Thus, we have shown how Equation 3 can be interpreted as a cross-entropy loss. Equation 3 with general $R(f(x),y)$ can be seen as a cross-entropy loss with a “smoothed” ground truth distribution $q_{x}$ given by $q_{x}\propto p_{W}[f]\cdot R(f(x),y).$
REINFORCE Loss
Reinforcement-learning losses are effective for exploring large search spaces. We use a modification of the REINFORCE [34] loss because it is relatively simple to implement. Future work could explore more sophisticated variants of this algorithm, such as Proximal Policy Optimization [46].
The standard REINFORCE loss applied to OccamLLM gives
$\mathcal{L}(x,y;W)=-\frac{1}{N}\sum_{f\sim p_{W}}R(f(x),y)\cdot\log p_{W}[f].$
Note that for sparse $R$, there will be very few nonzero $R(f(x),y)$ sampled, so, since we are dividing by $N$, the gradient signal will be small. We modify REINFORCE by dividing by the the sum of the rewards for all samples instead of by $N$ to ensure that correct functions sampled only a few times still receive a large training step update. This once again produces Equation 3.
We find that using a delta function for our reward is most effective because it most accurately represents the sparse reward of the problem. Further, as shown above, this loss provides a Monte-Carlo estimate of the the cross entropy loss. Due to the sparse reward, many samples may initially be required to obtain an accurate estimate of the loss. However, as OccamNet approaches the desired distribution, the loss’s sample efficiency will improve.
Appendix D Background on OccamNet
This section is heavily modified from [11].
@@ Appendix D Background on OccamNet
    In Section D.1, we describe OccamNet’s architecture in more detail.
@@
    In Section D.2, we describe OccamNet’s sampling process.
@@
    In Section D.3, we describe OccamNet’s probability distribution.
@@
    In Section D.4, we describe OccamNet’s initialization process.
D.1 OccamNet Architecture
As described in the main text, we start from a predefined collection of $N$ primitive functions $\mathcal{P}$. OccamNet represents a distribution over compositions of functions in $\mathcal{P}.$ From now on, we denote the $i$th primitive function in the $l$th layer as $\phi_{i}^{(l)}$. We begin indexing the primitives from 0 and the layers from 1, because we treat the inputs as the 0th layer. So, for example, in Figure 5a, $\phi_{2}^{(1)}=\phi_{2}^{(2)}=\phi_{2}^{(3)}=\mathsf{sin}$.
Figure 5: $a)$ An example OccamNet, with image layers boxed in green and arguments layers boxed in blue. We denote the inputs as the 0th image layer and the outputs as the $(L+1)$th arguments layer. Nodes in the arguments layers are represented with a P because of their probabilistic nature. $b)$ A demonstration of the dropped connections from sampled paths in OccamNet. All light grey paths are dropped from the final symbolic form of the sampled function because they are not directly connected to the outputs.
Each OccamNet layer consists of two sublayers, which we denote the arguments and image sublayers, shown in Figure 5a. For an $L$-layer OccamNet, each of these sublayers is reproduced $L$ times. The $l$th softmax layer connects the $(l-1)$th image layer with the $l$th arguments layer. For $1\leq l\leq L$, we denote the $l$th arguments sublayer hidden state as $\widetilde{\mathbf{h}}^{(l)}$ and the $l$th image sublayer hidden state as $\mathbf{h}^{(l)}$. So, $\widetilde{\mathbf{h}}^{(2)}$ would represent the middle layer of nodes labeled $P$ in Figure 5a. We further write
$\mathbf{\widetilde{h}}^{(l)}=\left[\widetilde{h}^{(l)}_{1},\dots,\widetilde{h}^{(l)}_{M^{(l)}}\right]^{\top},\;\;\mathbf{h}^{(l)}=\left[h^{(l)}_{1},\dots,h^{(l)}_{N^{(l)}}\right]^{\top},$ (9)
@@ D.1 OccamNet Architecture
$M^{(l)}=\sum_{0\leq k<N^{(l)}}\alpha\left[\phi_{k}^{(l)}\right],$
$N^{(l)}$ is the number of primitives in layer $l$, and $\alpha[\phi]$ is the arity of function $\phi.$ We also define $\mathbf{h}^{(0)}$ to be the input layer (an image sublayer) and $\mathbf{\widetilde{h}}^{(L+1)}$ to be the output layer (an arguments sublayer).
In a standard OccamNet layer, each primitive is repeated exactly once in each layer. However, in Complete OccamNet, each primitive in the $l$th layer is repeated $A^{L-l}$ times, where $A$ is the maximum arity of the primitives. This is shown in Figure 6 in the transition from 6a to 6b. Complete OccamNet also concatenates each image layer to the next image layer, as shown in Figure 6c.
D.2 Sampling from OccamNet
In this section, we more carefully describe OccamNet’s sampling process. We sample a connection to each arguments layer node from the distribution given by the softmax of the softmax-layer weights leading to that node. In particular, if $\mathbf{w}_{i}^{(l)}$ are the weights of the $l$th softmax layer leading to the $i$th node of the $l$th argument’s layer, when we sample we produce a sparse matrix
@@ D.2 Sampling from OccamNet
where the $\mathsf{SAMPLE}$ function samples a one-hot row vector for each row based on the categorical probability distribution defined by $\mathsf{softmax}(\mathbf{w})$. To evaluate this sample, we simply evaluate a forward pass through the network, treating the sampled sparse matrices from the softmax layers as the weights of linear layers:
@@
$h^{(l)}_{i}=\phi_{i}^{(l)}\left(\widetilde{h}_{j}^{(l)},\dots,\widetilde{h}_{j+\alpha[\phi_{i}^{(l)}]-1}^{(l)}\right),\;\;j=\sum_{0\leq k<i}\alpha\left[\phi_{k}^{(l)}\right].$ (12)
Figure 6: The progression of enhancements leading to a Complete OccamNet from a standard OccamNet. a) A standard OccamNet without repeated activations or skip connections. b) The same OccamNet as in a) with activations repeated in earlier layers. c) The same OccamNet as in b) with added skip connections. This is a Complete OccamNet.
D.3 OccamNet’s Probability Distribution
OccamNet parametrizes a probability distribution over all functions which it can sample. In particular, when OccamNet samples a function, it is really sampling a directed acyclic graph (DAG) which defines a computational path to compute a function. The probability of sampling a computational graph is equal to the product of the probabilities of the connections in the DAG which are connected to the output node.
Note that multiple computational graphs can correspond to the same function. In this paper, when we refer to a function sampled from OccamNet or the probability of a function according to OccamNet, we use function as a shorthand for a particular computational graph corresponding to that function. Although this underspecifies the computational graph in question, this is never an issue because we always refer to functions in abstract.
@@ D.3 OccamNet’s Probability Distribution
D.4 Initialization
This section describes how we calculate $\mathbf{W}^{*}$ from the main text. We wish to initialize $\mathbf{W}^{*}$ such that $p_{W^{*}}[f_{1}]=p_{W^{*}}[f_{2}]$ for all $f_{1}$ and $f_{2}$. Below, we assume that skip connections do not exist. However, the algorithm also works for skip connections, requiring only a small modification to Equation 14.
Unfortunately, such an initialization is impossible for any OccamNet with two or more layers containing primitives with more than one argument. However, it is possible to initialize OccamNet such that a lower bound $q_{W^{*}}$ of the true probability $p_{W^{*}}$ is independent of $f.$
Define the probability of a function $f$ up to a given node as the product of the probabilities of the edges that lead to that node in the DAG of $f$. Intuitively, $q_{W}[f]$ approximates $p_{W}[f]$ by maintaining a lower bound on the probability of $f$ up to each node of an OccamNet and propagating that lower bound through the computational graph given by $f$.
To define $q_{W}$ more precisely, let $q_{i}^{(l)}[f]$ and $\widetilde{q}_{i}^{(l)}[f]$ be the probability bounds corresponding to the $i$th node of the $l$th image or arguments sublayer. We have suppressed the dependence on $W$ for notational convenience. We compute these probabilities starting with the inputs, for which we set $q_{i}^{(0)}=1$. We then propagate probabilities to the arguments layers according to
$\widetilde{q}_{i}^{(l+1)}=\mathsf{softmax}(\mathbf{w}_{i}^{(l+1)})_{j}q_{j}^{(l)},$ (13)
where $j$ is the node in the $l$th image layer which $f$ connects to the $i$th node of $(l+1)$th arguments layer. Similarly, we propagate probabilities to the image layers according to
$q_{i}^{(l)}=\prod_{k=n}^{n+\alpha[\phi_{i}^{(l)}]-1}\widetilde{q}_{k}^{(l)},\hskip 10.0ptn=\sum_{j=1}^{i-1}\alpha\left[\phi_{j}^{(l)}\right].$ (14)
Finally, we define $q_{W}[f]=q_{0}^{(L+1)}[f].$
In practice $q_{W}[f]\leq p_{W}[f]$, where equality holds for many functions. In fact, $q_{W}[f]<p_{W}[f]$ only when part of the DAG of $f$ is used as input to two different arguments nodes. In cases such as these, the portion of the DAG that is used twice multiplicatively contributes the probability of its edges to $q_{W}[f]$ twice, artificially suppressing its value. However, because $q_{W}[f]$ is a lower bound, initializing $W^{*}$ to equalize $q_{W^{*}}$ still has the desired effect of ensuring adequate coverage for each $f$ in the initial probability distribution of OccamNet.
With this primer, we can now define the algorithm to initialize ${W^{*}}$ such that $q_{W^{*}}[f]$ is uniform.
The algorithm traverses through OccamNet layer by layer and establishes as an invariant that, after assigning the weights up to the $l$th layer, $\widetilde{q}_{i}^{(l)}[f]$ are equal for all $i$ and $f$. This implies that, after assigning the weights up to the $l$th layer, $q_{i}^{(l)}[f]$ are equal for all $f$, but not necessarily for all $i$. We denote the common value of $\widetilde{q}_{i}^{(l)}[f]$ as $\widetilde{q}^{(l)}$ and the common value of $q_{i}^{(l)}[f]$ as $q_{i}^{(l)}.$
The algorithm starts with input layer, where $q_{i}^{(0)}=1$ automatically. Once the invariant above is true for a given $l,$ the algorithm sets
$\left(\mathbf{w}_{i}^{*(l+1)}\right)_{j}=\log\left(\frac{\min_{k}\left(q_{k}^{(l)}\right)}{q_{j}^{(l)}}\right)$ (15)
for all $i,j$, where $\left(\mathbf{w}_{i}^{*(l+1)}\right)_{j}$ denotes the weight connecting the $j$th node in the $l$th image layer to the $i$th node in the $(l+1)$th arguments layer. This establishes the invariant for $l+1$ because
$q_{j}^{(l)}\mathsf{softmax}(\mathbf{w}_{i}^{*(l+1)})_{j}$ $=\frac{q_{j}^{(l)}\exp\left[\left(\mathbf{w}_{i}^{*(l+1)}\right)_{j}\right]}{\sum_{k}\exp\left[\left(\mathbf{w}_{i}^{*(l+1)}\right)_{k}\right]}$
$=\frac{q_{j}^{(l)}\min_{k}\left(q_{k}^{(l)}\right)/q_{j}^{(l)}}{\sum_{k}\min_{m}\left(q_{m}^{(l)}\right)/q_{k}^{(l)}}$
$=\frac{1}{\sum_{k}1/q_{k}^{(l)}},$
which is a constant over both $i$ and $j$, so $\widetilde{q}_{i}^{(l+1)}[f]$ is a constant over both $i$ and $f.$ The algorithm repeats the above procedure until it has traversed the entire network.